OpenStudy (anonymous):

x − 2y = 14 x + 3y = 9

OpenStudy (anonymous):

Solve for the system of equations

OpenStudy (uri):

These are two equations right?

OpenStudy (anonymous):

Yea

OpenStudy (uri):

@evilmath huh?

OpenStudy (anonymous):

Thank you

OpenStudy (uri):

that was wrong?

OpenStudy (uri):

Also why did you delete it @evilmath

OpenStudy (anonymous):

Yea i cant find it anymore

OpenStudy (uri):

That was weird.

OpenStudy (anonymous):

@uri can you help me?

OpenStudy (uri):

Yes I'll help you

OpenStudy (uri):

2x − 4y =28(equation 1) 2x + 6y =18(equation 2) - - - ======== -10y=10 y=10/-10 y=-1 Put the valur of y in equation 1 2x+4=28 2x=28-4 x=24/2 x=12

OpenStudy (uri):

to eliminate x we mulltiply both the equations by 2,Understand?

OpenStudy (anonymous):

A little thank you

OpenStudy (anonymous):

So the answer is (12,-1)?

OpenStudy (uri):

Yes :)